How many different possible committees of
7 men and 7 women
can be selected from a pool of candidates containing
9 men and 8 women?

Answer by Edwin McCravy(20054)

9 choose 7 times 8 choose 7 = C(9,7)·C(8,7) = 36·8 = 288.